### v2

@@ 6 Experiments
In this section, we outline the experimental setup used to evaluate the performance of the proposed MITST model for predicting BG levels in ICU patients. We then present our experimental results, organized around four key research questions: Q1: How does the model’s predictive performance compare to the baseline across different BG level classes (Section 6.2)? Q2: When the model makes mistakes, what is the clinical severity of these errors (Section 6.3)? Q3: How does the performance of the model degrade as the time to the next BG measurement increases (Section 6.4)? Q4: How does the model perform across different diabetic patient cohorts (Section 6.5)? The source code, along with all implementation details, is publicly available at: https://github.com/zavareh89/MITST.
@@ 6.2 Prediction quality comparison
When comparing the performance of the two models, MITST demonstrates superior capability, particularly in predicting hypoglycemia as shown in Fig. 3 and Table. 2. It achieves an AUROC of 0.915 (CI: 0.912–0.917), significantly outperforming RF ($p<0.001$), which reached 0.862 (CI: 0.858–0.865). Similarly, MITST’s AUPRC attains 0.247 (CI: 0.240–0.254), surpassing RF’s 0.208 (CI: 0.201–0.214; $p<0.001$). In the case of hyperglycemia, although the gap is narrower, MITST still maintains an advantage, with an AUROC of 0.909 (CI: 0.908–0.909) compared to 0.903 (CI: 0.902–0.903) for RF ($p<0.001$), and an AUPRC of 0.781 (CI: 0.779–0.783), slightly higher than RF’s 0.767 (CI: 0.765–0.769; $p<0.001$). The lower AUPRC values observed for the minority classes—hypoglycemia and hyperglycemia—are an anticipated outcome of the class imbalance present in the dataset. Since precision-recall curves are more sensitive to class distribution, the smaller number of positive examples in these categories naturally results in lower AUPRC values. Despite this, MITST’s relatively higher AUPRC for these minority classes highlights its robustness in handling imbalanced data. These results demonstrate that the proposed MITST model not only improves predictive performance for minority classes, but also maintains competitive performance for the majority class.
@@ 7 Discussion
In this study, we have developed and validated a novel framework, MITST, for predicting BG levels in ICU patients using a multi-source irregular time-series dataset. Leveraging Transformers, MITST effectively handles the complexity and irregularity of EHR data at multiple levels of granularity, outperforming traditional approaches, particularly in predicting critical conditions such as hypoglycemia and hyperglycemia. MITST showed 1.7% improvement in macro-averaged AUROC ($p<0.001$) and 1.8% increase in macro-averaged AUPRC ($p<0.001$) compared to a state-of-the-art random forest model, significantly enhancing prediction accuracy for at-risk glucose levels. These results emphasize the model’s capacity to provide timely, accurate predictions, enabling clinicians to take proactive measures that improve patient outcomes. Additionally, MITST maintains consistent performance across specific clinical contexts, such as diabetic cohorts. As shown in Table. 3, the model demonstrated robustness in predicting glucose fluctuations for both type 1 and type 2 diabetes patients. This consistency highlights MITST’s flexibility and scalability by integrating diverse data sources, offering a reliable solution for real-time monitoring and decision support across diverse patient populations in ICU settings, helping mitigate the risk of dysglycemia.
@@
While MITST offers substantial improvements in predicting critical glucose levels, certain limitations must be considered. The complexity of Transformers, which involve multiple layers of self-attention mechanisms, results in longer training times due to the need for large-scale data and extensive computations during model optimization. However, once trained, the model’s inference time is highly efficient—on the order of a few milliseconds—making it well-suited for real-time clinical tasks. The fast inference time mitigates concerns about computational efficiency during real-time deployment, allowing the model to integrate seamlessly into clinical workflows. Although the initial training process is resource-intensive, it is a one-time cost that does not affect the model’s performance or scalability in real-time environments.
@@ 8 Conclusions and future work
While the current study demonstrates the effectiveness of supervised learning in predicting BG levels, future work could explore the application of self-supervised learning techniques. This approach would be particularly useful in scenarios where labeled data is scarce or difficult to obtain. Additionally, future work could explore the use of multi-task learning [45] to further enhance MITST’s versatility. Multi-task learning would enable the model to simultaneously learn multiple related tasks—such as predicting other vital signs or clinical outcomes—by sharing representations across tasks, improving generalization and efficiency. Another direction for future work is the application of large language models (LLMs) in EHR data analysis, utilizing their capability to extract insights from unstructured clinical notes and improve multi-source data integration. Also, from an ethical standpoint, it is essential to recognize potential biases within the model due to the underlying dataset, which may not fully represent all ICU patient populations. To mitigate this risk, future studies should prioritize fairness and explainability [46], ensuring that the model does not disproportionately benefit or harm any specific subgroup of patients.

### v3

@@ 6 Experiments
In this section, we outline the experimental setup used to evaluate the performance of the proposed MITST model for predicting BG levels in ICU patients. We then present our experimental results, organized around five key research questions: Q1: How does the model’s predictive performance compare to the baseline across different BG level classes (Section 6.2)? Q2: When the model makes mistakes, what is the clinical severity of these errors (Section 6.3)? Q3: How does the performance of the model degrade as the time to the next BG measurement increases (Section 6.4)? Q4: How does the model perform across different diabetic patient cohorts (Section 6.5)? Q5: Can the model generalize to a different clinical prediction task, such as in-hospital mortality (Section 6.6)? The source code, along with all implementation details, is publicly available at: https://github.com/zavareh89/MITST.
@@ 6.1.1 Competing method
To further strengthen our comparative analysis, we employed STraTS [30]—a recent Transformer-based model for irregular clinical time series—as an additional baseline selected from the deep learning-based methods. We carefully adapted the original open-source codebase for the BG prediction task, extracting 102 features from our dataset and formatting them as required triplets (feature name, value, timestamp). Due to STraTS’s architectural constraints, only features with continuous values could be incorporated, as its initial embedding layer supports only continuous-valued inputs. This limitation prevented the inclusion of categorical sources such as diagnoses or treatments, thereby reducing the richness of input data. For fairness, we employed the same patient splits, data preprocessing, and class imbalance handling as in our method. The maximum sequence length, which was 880 in the original STraTS implementation to prevent memory overflow, was increased to 1024 observations in our adaptation to better accommodate the high temporal density of our data. Hyperparameters were tuned for our task to ensure optimal performance.
@@ 6.1.3 Evaluation metrics
6.1.4 Class imbalance handling
To assess the effect of class imbalance handling on the baseline, we implemented a balanced RF model [43] as a comparative variant. In this approach, each decision tree in the RF ensemble draws an equal number of bootstrap samples from both the minority and majority classes, effectively simulating a per-tree undersampling strategy. This configuration differs from the standard RF model used in the original study [18], which did not report any explicit balancing method. As shown in Table. A3 in the appendix, the balanced RF modestly improved AUROC and sensitivity for the hypoglycemia class but degraded overall macro-averaged performance for these two metrics. Hyperparameters for both RF variants were tuned via cross-validation on the validation set. Based on these findings, we retained the original unbalanced RF setup as the primary baseline in our main results to ensure consistency with prior work and to present the most competitive benchmark.
@@ 6.1.5 Experimental parameter and training settings
To further examine the robustness of our evaluation, we also conducted a 5-fold patient-level cross-validation using the same features, model configuration, and training procedures. As shown in Table. A4 in the appendix, the average performance across folds remained consistent with the original hold-out results—for instance, AUROC and sensitivity for hypoglycemia were 0.917 $\pm$ 0.001 and 0.837 $\pm$ 0.002, closely matching the original split values (0.915 and 0.841, respectively). These results confirm the stability and generalizability of the model’s performance across patient subsets. Furthermore, given that the dataset includes approximately 2.6 million training and 0.74 million test examples, the original hold-out split remains clinically realistic. It mimics prospective deployment scenarios, where the model is applied to entirely unseen cohorts.
@@ 6.2 Prediction quality comparison
When comparing the performance of the two models, MITST demonstrates superior capability, particularly in predicting hypoglycemia—a clinically critical but rare event—raising AUROC from 0.862 (CI: 0.858–0.865) to 0.915 (CI: 0.912–0.917) (+5.3 pp) and sensitivity from 0.769 (CI: 0.76.5–0.774) to 0.841 (CI: 0.83.8–0.84.5) (+7.2 pp) as shown in Fig. 3 and Table. 2. Our test set includes over 14,000 hypoglycemia cases, so an improvement of this magnitude would result in roughly 1,000 additional early detections. In critical care settings, where every minute counts, this could enable a significant clinical improvement of timely interventions, potentially reducing both morbidity and mortality. Similarly, MITST’s AUPRC attains 0.247 (CI: 0.240–0.254), surpassing RF’s 0.208 (CI: 0.201–0.214; $p<0.001$). In the case of hyperglycemia, although the gap is narrower, MITST still maintains an advantage, with an AUROC of 0.909 (CI: 0.908–0.909) compared to 0.903 (CI: 0.902–0.903) for RF ($p<0.001$), and an AUPRC of 0.781 (CI: 0.779–0.783), slightly higher than RF’s 0.767 (CI: 0.765–0.769; $p<0.001$). The lower AUPRC values observed for the minority classes—hypoglycemia and hyperglycemia—are an anticipated outcome of the class imbalance present in the dataset. Since precision-recall curves are more sensitive to class distribution, the smaller number of positive examples in these categories naturally results in lower AUPRC values. Despite this, MITST’s relatively higher AUPRC for these minority classes highlights its robustness in handling imbalanced data. These results demonstrate that the proposed MITST model not only improves predictive performance for minority classes, but also maintains competitive performance for the majority class.
@@
To compare MITST with more advanced deep learning models, we implemented and evaluated STraTS [30], a recent Transformer-based architecture specifically designed for irregular clinical time series. The results are presented in Table. 3, using the same settings as in Table. 2. While STraTS represents a conceptually powerful architecture, its performance in our setting was notably inferior to RF and MITST models. On average, STraTS underperformed even compared to the RF baseline, with a drop of approximately 3.5 pp in AUROC and 5.4 pp in AUPRC across all classes. We attribute this discrepancy to several limitations of STraTS when applied to our task. First, STraTS only supports continuous-valued features and lacks native support for categorical inputs such as diagnosis codes or treatment types, which are critical components of our model’s input space. As a result, a small but clinically informative portion of features was excluded from its representation. Second, its original design and hyperparameter choices were tailored for the simpler task of in-hospital mortality prediction, which involves a coarser temporal resolution. In contrast, BG prediction is more temporally sensitive and clinically nuanced. Despite tuning STraTS for our use case, it remained less effective at capturing the fine-grained, multi-source dependencies essential for accurate BG forecasting.
@@
Table 3: Performance metrics comparison across different classes (euglycemia, hypoglycemia, hyperglycemia) and their macro-average values for STraTS [30] and MITST models.
STraTS [30] MITST model (ours)
hypo hyper euglycemia macro avg hypo hyper euglycemia macro avg
prevalence 0.019 0.232 0.749 0.019 0.232 0.749
Metrics
AUROC 0.872 0.845 0.827 0.848 0.915 0.909 0.876 0.900
AUPRC 0.170 0.663 0.931 0.588 0.247 0.781 0.951 0.660
sensitivity 0.780 0.767 0.728 0.758 0.841 0.833 0.778 0.817
specificity 0.796 0.745 0.757 0.766 0.845 0.830 0.814 0.830
AUROC = Area Under the Receiver Operating Characteristic Curve; AUPRC = Area Under the Precision-Recall Curve.
@@ 6.5 Performance analysis for diabetic cohorts
6.6 Generalization to mortality prediction task
To assess the generalizability of MITST across tasks, we conducted an additional experiment on in-hospital mortality prediction [44]. This setting differs substantially from our original BG prediction task and serves as a strong test for model transferability. We first filtered the dataset to include only ICU stays longer than 24 hours that also contained recorded APACHE IVa scores (described below) [45], resulting in a total of 81,184 ICU stays. We used the same train/validation/test split as in the original task, consisting of 56,842, 8,035, and 16,307 stays, respectively. The average in-hospital mortality rate in this subset was 10.1%. We used the pretrained MITST model from the BG prediction task and fine-tuned it on this new task. To balance transferability and efficiency, we froze all weights except for the fusion layer and the final Transformer module (responsible for semantic multi-source representation). This strategy allows the model to retain the source-specific temporal encoding learned from the original task, while adapting higher-level semantic integration to the new target. We found this configuration to offer a good trade-off between retaining useful representations and allowing task-specific adaptation.
For comparison, we implemented STraTS on this task using the same features and triplet-based input structure as previous task. Unlike MITST, STraTS was trained from scratch rather than fine-tuned from its BG-trained model, due to its poor performance on that task, as shown in Section 6.2. We used the same hyperparameter settings as those reported in the original STraTS paper for in-hospital mortality prediction. As an additional baseline, we included the APACHE IVa score [45], a well-established illness severity scoring system widely used in ICUs for outcome prediction. APACHE IVa incorporates a combination of physiological variables, chronic health conditions, and admission characteristics to estimate in-hospital mortality risk. As discussed in [44], such scoring systems have shown strong predictive performance and remain the clinical standard in many ICUs. It is important to note, however, that APACHE IVa is a proprietary system, and the exact implementation details (e.g., variable mappings, coefficients) are not publicly available. For this reason, we used the APACHE IVa scores directly from the eICU database rather than attempting to replicate the scoring algorithm. Including APACHE IVa as a baseline allows us to benchmark our model against a domain-validated, interpretable alternative.
The comparative results for MITST, STraTS, and APACHE IVa are presented in Table. 5, using AUROC and AUPRC as the primary evaluation metrics. Despite being fine-tuned on the mortality prediction task with only a subset of its parameters updated (i.e., the fusion layer and final Transformer module), MITST achieved performance comparable to STraTS, which was trained from scratch on this task. This suggests that the source-specific temporal representations learned during pretraining on BG prediction generalized effectively to a different clinical outcome. We note that no additional task-specific hyperparameter optimization was applied to MITST; therefore, its performance could potentially be further improved through targeted tuning. These findings demonstrate the transferability and adaptability of MITST across prediction tasks. In contrast, APACHE IVa outperformed both deep learning models, which can likely be attributed to its highly curated and task-specific input space, optimized specifically for mortality risk estimation in ICUs. By comparison, the feature set used to train MITST and STraTS was originally selected for BG prediction and was intended to be general-purpose. It also lacked several input variables known to be relevant to mortality risk (e.g., detailed ventilator settings, additional lab measurements). Future work could address this limitation by expanding the input space to include a broader range of sources and richer clinical signals per source, potentially narrowing the gap with expert-designed scoring systems.
Table 5: Comparison of AUROC and AUPRC for in-hospital mortality prediction using three models: APACHE IVa (clinical scoring system) [45], STraTS (deep learning baseline) [30], and MITST (fine-tuned from BG prediction task).
Metric APACHE IVa [45] STraTS [30] MITST (fine-tuned)
AUROC 0.824 0.804 0.806
AUPRC 0.410 0.382 0.379
AUROC = Area Under the Receiver Operating Characteristic Curve; AUPRC = Area Under the Precision-Recall Curve.
@@ 7 Discussion
In this study, we have developed and validated a novel framework, MITST, for predicting BG levels in ICU patients using a multi-source irregular time-series dataset. Leveraging Transformers, MITST effectively handles the complexity and irregularity of EHR data at multiple levels of granularity, outperforming traditional approaches, particularly in predicting critical conditions such as hypoglycemia and hyperglycemia. MITST showed 1.7 pp improvement in macro-averaged AUROC ($p<0.001$) and 1.8 pp increase in macro-averaged AUPRC ($p<0.001$) compared to a state-of-the-art random forest model, significantly enhancing prediction accuracy for at-risk glucose levels. These results emphasize the model’s capacity to provide timely, accurate predictions, enabling clinicians to take proactive measures that improve patient outcomes. Additionally, MITST maintains consistent performance across specific clinical contexts, such as diabetic cohorts. As shown in Table. 4, the model demonstrated robustness in predicting glucose fluctuations for both type 1 and type 2 diabetes patients. This consistency highlights MITST’s flexibility and scalability by integrating diverse data sources, offering a reliable solution for real-time monitoring and decision support across diverse patient populations in ICU settings, helping mitigate the risk of dysglycemia.
@@
While MITST offers substantial improvements in predicting critical glucose levels, certain limitations must be considered. The complexity of Transformers, which involve multiple layers of self-attention mechanisms, results in longer training times due to the need for large-scale data and extensive computations during model optimization. However, once trained, the model’s inference time is highly efficient—on the order of a few milliseconds—making it well-suited for real-time clinical tasks. The fast inference time mitigates concerns about computational efficiency during real-time deployment, allowing the model to integrate seamlessly into clinical workflows. Although the initial training process is resource-intensive, it is a one-time cost that does not affect the model’s performance or scalability in real-time environments and can be justified by its meaningful impact in detecting rare yet critical conditions like hypoglycemia. Standard models often overlook such infrequent classes due to skewed class distributions, whereas MITST maintains high sensitivity through its end-to-end, attention-based integration of diverse clinical data.
@@
Translating predictive models such as MITST into clinical practice demands more than high performance; it requires addressing implementation feasibility, user interaction, and organizational readiness. To demonstrate this, we developed a proof-of-concept web interface that enables real-time interaction with MITST. Users can adjust input features and immediately observe resulting prediction changes, thereby simulating its potential role in a clinical decision support tool (see Appendix).
In typical ICU settings, essential inputs (e.g., heart rate, respiratory rate) are automatically captured by patient monitors and EHR systems, with only a few variables (like diagnoses or medications) needing manual entry—though these, too, can be integrated with EHRs. While the current prototype emphasizes the prediction flow, a complete system would require seamless integration with existing clinical infrastructure to minimize user burden. Moreover, developing a comprehensive Clinical Decision Support System (CDSS) entails addressing organizational, technical, and regulatory challenges—including workflow integration, coordination among clinicians, IT staff, and administrators, along with patient safety evaluations and usability testing [47]. Thus, while promising, successful clinical translation remains a significant future undertaking.
@@ 8 Conclusions and future work
While the current study demonstrates the effectiveness of supervised learning in predicting BG levels, future work could explore the application of self-supervised learning techniques. This approach would be particularly useful in scenarios where labeled data is scarce or difficult to obtain. Additionally, future work could explore the use of multi-task learning [48] to further enhance MITST’s versatility. Multi-task learning would enable the model to simultaneously learn multiple related tasks—such as predicting other vital signs or clinical outcomes—by sharing representations across tasks, improving generalization and efficiency. Another direction for future work is the application of large language models (LLMs) in EHR data analysis, utilizing their capability to extract insights from unstructured clinical notes and improve multi-source data integration. Also, from an ethical standpoint, it is essential to recognize potential biases within the model due to the underlying dataset, which may not fully represent all ICU patient populations. To mitigate this risk, future studies should prioritize fairness and explainability [49], ensuring that the model does not disproportionately benefit or harm any specific subgroup of patients. Finally, a proof-of-concept user interface is developed to demonstrate potential deployment; however, translating this model into a full Clinical Decision Support System (CDSS) would require significant clinical and organizational integration.
@@ 17 Temperature nurseCharting location value
Table A3: Comparison of performance metrics for the original and balanced Random Forest (RF) models across glycemic classes (hypoglycemia, hyperglycemia, euglycemia), including macro-averaged values. The original RF model [18] does not apply any explicit class balancing strategy, whereas the balanced RF variant uses per-tree class-balanced bootstrap sampling—drawing equal numbers of samples from both minority and majority classes in each decision tree. This approach is analogous to the undersampling technique used in our deep model (MITST). While the balanced RF slightly improves AUROC and sensitivity for the hypoglycemia class, it results in decreased macro-averaged performance across most metrics. Hyperparameters for both RF models were tuned via cross-validation on the validation set.
Original RF model [18] Balanced RF model
hypo hyper euglycemia macro avg hypo hyper euglycemia macro avg
prevalence 0.019 0.232 0.749 0.019 0.232 0.749
Metrics
AUROC 0.862 0.903 0.884 0.883 0.872 0.898 0.866 0.878
AUPRC 0.208 0.767 0.951 0.642 0.204 0.746 0.946 0.632
PPV 0.081 0.602 0.922 0.535 0.090 0.589 0.921 0.533
NPV 0.995 0.938 0.591 0.841 0.995 0.940 0.548 0.828
sensitivity 0.769 0.818 0.816 0.801 0.778 0.824 0.779 0.794
specificity 0.829 0.837 0.795 0.820 0.847 0.826 0.801 0.825
Table A4: Comparison of MITST model performance using (left) the original patient-level train/validation/test split from the main paper and (right) a 5-fold patient-level cross-validation. Metrics are reported per glycemic class (hypoglycemia, hyperglycemia, euglycemia) and as macro-averages. In the cross-validation setting, the reported values represent the mean and standard deviation across folds. Note that for cross-validation, we used a different random partitioning of patients than the original hold-out split used. The results demonstrate strong agreement between the two evaluation strategies.
MITST (original train/val/test split used in main paper) MITST (5-fold patient-level cross-validation)
hypo hyper euglycemia macro avg hypo hyper euglycemia macro avg
prevalence 0.019 0.232 0.749 0.019 0.232 0.749
Metrics
AUROC 0.915 0.909 0.876 0.900 0.917 $\pm$ 0.001 0.909 $\pm$ 0.001 0.878 $\pm$ 0.001 0.901 $\pm$ 0.001
AUPRC 0.247 0.781 0.951 0.660 0.251 $\pm$ 0.003 0.780 $\pm$ 0.001 0.952 $\pm$ 0.001 0.661 $\pm$ 0.001
PPV 0.096 0.596 0.926 0.539 0.098 $\pm$ 0.001 0.592 $\pm$ 0.002 0.927 $\pm$ 0.001 0.539 $\pm$ 0.001
NPV 0.996 0.943 0.551 0.830 0.996 $\pm$ 0.001 0.944 $\pm$ 0.001 0.550 $\pm$ 0.005 0.830 $\pm$ 0.002
sensitivity 0.841 0.833 0.778 0.817 0.837 $\pm$ 0.002 0.834 $\pm$ 0.001 0.779 $\pm$ 0.004 0.817 $\pm$ 0.002
specificity 0.845 0.830 0.814 0.830 0.852 $\pm$ 0.002 0.828 $\pm$ 0.002 0.814 $\pm$ 0.003 0.831 $\pm$ 0.001
Supplementary Material: Interactive MITST Demo System
We developed a web-based proof-of-concept interface that executes the pre-trained MITST model on custom input data. Hosted on an Aalborg University server, this interface simulates clinical inference for a hypothetical ICU patient. Note that it is intended solely for demonstration purposes and not for production deployment. A live demo is available at http://130.225.37.30:8501/.
System Overview and Usage
Figure S1 illustrates the main interface of the system, which is organized into multiple tabs representing different data sources. After entering or loading the required data, users can press the “Run MITST Model” button to obtain a predicted glucose level label (i.e., hypoglycemia, euglycemia, or hyperglycemia). The inference is performed on the server using a pretrained model.
To simplify usage, we provide six pre-filled example templates from the test set of the eICU database, accessible through dedicated buttons in the interface (see Figure S1). These include:
  • 3 examples for hypoglycemia: True Positive, False Positive, False Negative.
  • 3 examples for hyperglycemia: True Positive, False Positive, False Negative.
Selecting any of these automatically fills all input tabs with the corresponding data (Figure S3). Users can modify any value and re-run the model to observe how predictions change in real-time.
Input Structure and Tabs
The system handles heterogeneous clinical data following the MITST training format. Input is divided across multiple tabs—each corresponding to one or more data sources—to streamline user interaction and visualization.
  • Vital Signs: High-frequency signals like HR, RR, SpO2, blood pressures. Users enter comma-separated numeric values with corresponding timestamp offsets (see Figure S2).
  • Medications: Discrete entries for medications and infusions, including categorical fields (e.g., route, frequency) selected via drop-downs, and numeric dosage values entered via sliders.
  • Lab Results: Similar to medications, lab name and result are entered, along with timestamp offset.
  • Scores: Includes GCS and sedation scores. Each record consists of a categorical score and offset.
  • IO and IO Total: IO tab supports both discrete and total values. Total intake/output values are entered similarly to vital signs, while IO cell types (e.g., dialysis) are entered individually.
  • Past History and Treatment: Categorical variables like comorbidities and treatments are entered one by one with drop-downs.
  • Diagnosis: Supports a broad set of diagnoses and priorities, and also separate “admission diagnosis” (addx).
  • Demographics and Hospital Info: Includes static features like weight, height, gender, age, and hospital and unit metadata.
Several tabs include multiple data sources grouped logically (e.g., the “Vital Signs” tab includes HR, RR, SpO2, BP measurements). This design improves usability and visualization.
Notes on Limitations and Implementation
  • Min and Max Limits: All numeric sliders (e.g., for dosage or lab values) are bounded using thresholds based on the mean $\mu$ and the standard deviation $\sigma$ calculated from training data. These are not clinically validated limits and are used purely for demonstration.
  • Manual Data Entry: In a real-world scenario, clinicians and nurses would not manually input most of the fields. Many values such as vitals and labs are collected automatically through hospital systems and monitoring equipment. Only specific fields (e.g., diagnosis or treatments) may require human input.
  • Efficiency: The current implementation is not optimized for performance and is intended for proof-of-concept only. Future work may focus on inference time improvements and front-end optimizations.
  • Deployment: The model is pretrained offline and deployed for inference only. All processing happens server-side using CPU.
Figure S1: Main interface of the MITST demo system. Users can choose pre-filled examples or manually navigate between tabs to provide inputs. Pressing “Run MITST Model” button triggers inference.
Figure S2: Vital signs tab. Users enter comma-separated values and time offsets for each signal, reflecting irregular time-series format of real ICU data.
Figure S3: Example of a pre-filled case for hypoglycemia (False Positive). Users can modify any value to test sensitivity of the model to input changes.